OpenStudy (anonymous):

Solve by graphing. (If the solution is inconsistent, enter INCONSISTENT. If the solution is dependent, enter DEPENDENT.) x + y=2 x – y =6

OpenStudy (anonymous):

wouldnt this be inconsistent

OpenStudy (anonymous):

what does inconsistent mean? and what does dependent mean?

OpenStudy (anonymous):

when you graph the lies and they are the same they are dependent

OpenStudy (anonymous):

lines i mean

OpenStudy (anonymous):

well these aren't that, obviously

OpenStudy (anonymous):

if the lines are parallel and have no solution then the are inconsistent

OpenStudy (anonymous):

the lines aren't parallel

OpenStudy (anonymous):

have you graphed these?

OpenStudy (anonymous):

so it will be consistent

OpenStudy (anonymous):

no but if you will turn this to slope intercept form, you'll come up with the same conclusion

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

move the x to the other side, and be careful with the signs

OpenStudy (anonymous):

so what would x= and what would y=

OpenStudy (anonymous):

you want to get y = Mx + B

OpenStudy (anonymous):

for each one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the first line x+y=2: y=-x+2 has a slope of -1. The second line x-y=6 x-y=6 -y=-x+6 y=x-6 has a slope 1. Since they don't have the same slope, they are not parallel, so they intersect at one point. That point will be the solution to the system therefore the system is consistent

OpenStudy (anonymous):

ok I see what your saying

OpenStudy (anonymous):

nice question btw :D